OpenStudy (anonymous):

In triangle ABC, angle A is a right angle, and angle B is 45 degrees. Find BC if your answer is not an integer leave it in simplest form. I'll draw a pic and show the math I've done so far.

OpenStudy (anonymous):



OpenStudy (anonymous):

I know that since its a 45-90-45 degree angle I need to use the equation hypotenuse= radical 2 times the leg so I got x= radical 2 * 12

OpenStudy (jack1):

is angle b is 45 then angle c is 45 degrees also

OpenStudy (anonymous):

So then I got x=radical 24 simplified to 2 radical 6 but that isnt one of my options.

OpenStudy (jack1):

sqrt 288

OpenStudy (anonymous):

How did you get that

OpenStudy (jack1):

h^2 = a^2 + b^2 h^2 = a^2 + a^2 (isoceles triangle, 2x 45 degree angles and a right angle) h^2 = 2(a^2) and from your drawing a = 12 ft h^2 = 2 x (12^2) h^2 = 2 x 144 h^2 = 288 h = sqrt 288 = length B to C

OpenStudy (anonymous):

Oh okay. Thank you!

OpenStudy (jack1):

welcome